If 20 people complete the work in 30 days, then after how many days should 5 people go on leave so that the work is completed in 35 days.


answer is 5 days..
but how please share the complete methods to solve the question.
thanks

Answer by richwmiller(17219)

20 people working 30 days are 600 man work days to finish the job.
If it were 5 days it would be
20*5=100 man work days.
15 people working for 30 more days
15 * 30=450 man work days
100+450=550
that only adds up to 550 man work days which isn't enough to finish the job.
The correct answer is
20x+15y=600,
x+y=35
x = 15, y = 20
5 people leave after 15 days
and 15 people finish the work in 20 more days